# EEG-Based Consumer Behavior Prediction: An Exploration from Classical Machine Learning to Graph Neural Networks

### v1

@@ 6 Conclusion
In this study, the performance of different machine learning pipelines for predicting consumer behavior based on EEG signals in the field of neuromarketing was investigated. Two main approaches were compared: (1) classical and hybrid pipelines that rely on EEG statistical features and use conventional classifiers, and (2) graph-based approaches that exploit spectral features and brain connectivity representations and are modeled with the help of graph neural networks (GNNs). The experimental results showed that although classical and hybrid models have an acceptable performance in terms of accuracy, GNN-based approaches offer better performance for predicting the minor class.
The findings of this study demonstrate the significant potential of GNNs for modeling consumer behavior based on EEG and highlight their advantage in interactions of brain networks. However, classical methods remain competitive in situations where computational simplicity and interpretability are more important. This suggests that the choice of method depends on the specific circumstances and needs of each problem.
Despite these promising results, some limitations should be noted. This study was conducted on a limited dataset, which may affect the robustness and generalizability of the results. Also, only a limited set of connectivity features was examined, which allows for extension to richer multimodal data and the use of more advanced connectivity estimators in future research.
Overall, this study provides valuable empirical evidence confirming the effectiveness of graph-based learning in neuromarketing applications and introduces a way for the development of more accurate, scalable, and interpretable models for predicting consumer decisions. In the future, this line of research could be pursued by combining classic and deep learning models, developing real-time prediction frameworks, and integrating EEG with other modalities such as eye tracking or physiological signals. Such approaches could further advance neuromarketing applications towards practical implementation in real-world settings, while having a solid scientific basis.

### v2

@@ 7 Conclusion
In this study, different machine learning methods and pipelines have been utilized for predicting consumer behavior by using an EEG-based dataset with neuromarketing stimuli. Two main methods have been implemented and compared: (1) classical and ensemble models with three different pipelines that used statistical EEG features, and (2) graph-based models that used spectral features with brain connectivity representations and different graph neural network architectures. The results showed that, while the classical models and pipeline have a better performance in terms of accuracy, the GNNs perform better in predicting the minor class.
The findings of this study introduced the remarkable potential of graph neural networks in the EEG-based tasks by modelling the brain connectivity, which shows the interactions between the nodes. Due to the computational simplicity and interpretability, classical models can remain competitive in tasks where these measures are important, and this shows that the method selection can depend on the needs of a task and the problem specifications.
In spite of the significant results and methods that have been used in this study, the limitations should be mentioned. The study was conducted by using a limited dataset, and due to this, the results may affect the model in terms of robustness and generalizability. Furthermore, in future studies, different connectivity features can be examined to develop richer models and more accurate predictions.
Overall, this study provides a deeper insight into the effectiveness of graph-based approaches in neuromarketing applications, compares classical, ensemble, and GNN models from a neuromarketing perspective, and shows the applications of machine learning and deep learning methods in EEG-based consumer behavior predictions. The future research can be pursued by combining classic and deep learning models, using a wide range of datasets, and developing more robust models, utilizing different methods for construction on the brain connectivity graphs, and integrating EEG with other modalities such as eye tracking, physiological data, fMRI, or fNIRS data. These approaches can cause advancements in the field of neuromarketing in both practical insights and scientific prospective.